OpenStudy (anonymous):

Will fan and medal if you can help me with the answer and explain how to get it What is 2 2/3- 3/5?

OpenStudy (anonymous):

first, convert 2 and 2/3 to an improper fraction

OpenStudy (michele_laino):

hint: \[2\frac{2}{3} = 2 + \frac{2}{3} = \frac{8}{3}\]

OpenStudy (anonymous):

so take 3 times 2 and add 2, then put it over 3

OpenStudy (anonymous):

8/3?

OpenStudy (anonymous):

right!

OpenStudy (anonymous):

so subtract 3 from 8

OpenStudy (anonymous):

5

OpenStudy (anonymous):

and put that over 2 again, so your final answer is 5/2

OpenStudy (anonymous):

thanks for the help!

OpenStudy (anonymous):

no problem! :)

OpenStudy (michele_laino):

sorry, I think that we have: \[\frac{8}{3} - \frac{3}{5} = \frac{{8 \times 5 - 3 \times 3}}{{15}} = \frac{{31}}{{15}}\]

OpenStudy (anonymous):

oh my gosh that was really stupid of me! I completely forgot you need a common denominator! Forget the answer we got and do this ^^

OpenStudy (michele_laino):

please tell to @Ash_091 the right answer you are welcome! @jessmitz